University Hero is considering expanding operations beyond its healthy sandwiches. Jim Axelrod, vice president of marketing, would like to add a line of smoothies with a similar health emphasis. Each smoothie would include two free health supplements such as vitamins, antioxidants, and protein. Jim believes smoothie sales should help fill the slow mid-afternoon period. Adding the line of smoothies would require purchasing additional freezer space, machinery, and equipment. Jim provides the following projections of net sales, net income, and average total assets in support of his proposal.

Sandwiches Sandwiches
Only and Smoothies
Net sales $ 900,000 $ 1,500,000
Net income 170,000 260,000
Average total assets 500,000 900,000

Required:

1. Calculate University Hero’s return on assets, profit margin, and asset turnover for sandwiches only.
  

2. Calculate University Hero’s return on assets, profit margin, and asset turnover for sandwiches and smoothies.
  

3. Based on these ratios, what recommendation would you make if University Hero is considering expanding operations beyond its healthy sandwiches?
Answer #1

1) Calculate following :

Sandwiches only
Return on total assets 170000/500000 = 34%
Profit 170000/900000 = 18.89%
Asset turnover 900000/500000 = 1.8

2) Calculate following for sandwiches and smoothies

Return on assets 260000/900000 = 28.89%
Profit margin 260000/1500000 = 17.33%
Asset turnover 1500000/900000 = 1.67

3)No, Company should not expand their business they should sale only sandwiches